# A Multi-Fidelity Control Variate Approach for Policy Gradient Estimation

### v3

@@ 4 Multi-Fidelity Policy Gradients
Informally, when sampling trajectories from $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$, there are six sources of stochasticity: the stochasticity introduced by the policy $\pi_{\theta}$, that introduced by the initial state distribution $\Delta_{s_{I}}$, the stochasticity introduced by the high-fidelity and low-fidelity transition dynamics $p^{h}$ and $p^{l}$, and finally the stochasticity introduced by the two reward functions $R^{h}$ and $R^{l}$. Note that the stochasticity introduced by $\pi_{\theta}$ is implemented by the agent, and is thus under the control of the algorithm in the sense that the same random outcome $\omega$ may be used to generate actions from $\pi_{\theta}$ in both $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$. Similarly, the algorithm may fix the initial states in the low-fidelity environment to match those observed from the high-fidelity samples. On the other hand, the transition dynamics and environment rewards are generated by independent sources of stochasticity in the different environments. We accordingly define the outcome set of the probability space as $\Omega=\Omega_{\Delta_{s_{I}}}\times\Omega_{\pi}\times\Omega_{p^{h}}\times\Omega_{p^{l}}\times\Omega_{R^{h}}\times\Omega_{R^{l}}$. Here, each outcome $\omega^{\Delta_{s_{I}}}\in\Omega_{\Delta_{s_{I}}}$ defines a particular shared initial state. Meanwhile, each policy outcome $\omega^{\pi}\in\Omega_{\pi}$ corresponds to a sequence $\omega^{\pi}=\omega^{\pi}_{1},\omega_{2}^{\pi},\ldots,\omega_{T}^{\pi}$ that dictates the random sequence of actions selected by the policy $\pi_{\theta}$ in both environments. Conceptually, given the outcome $\omega_{t}^{\pi}$ at any timestep $t$ of a trajectory, the policy should deterministically output action $a^{h}_{t}=\pi_{\theta}(s_{t}^{h},\omega^{\pi})$ in the high-fidelity environment, and action $a^{l}_{t}=\pi_{\theta}(s_{t}^{l},\omega^{\pi})$ in the low-fidelity environment. Note that this does not necessarily imply that $a^{h}_{t}=a^{l}_{t}$, due to a potential difference in states $s^{h}_{t}$ and $s^{l}_{t}$ that stems from the different dynamics. Similarly, each outcome $\omega^{p_{h}}\in\Omega_{p^{h}}$ is a sequence $\omega^{p_{h}}=\omega_{0}^{p_{h}},\omega_{1}^{p_{h}},\ldots,\omega_{T}^{p_{h}}$ dictating the outcomes $s_{t+1}=p^{h}(s^{h}_{t},a^{h}_{t},\omega^{p_{h}})$ of the stochastic transitions in the high-fidelity environment. However, unlike the policy outcome sequence $\omega^{\pi}\in\Omega^{\pi}$, the transition outcome sequences $\omega^{p_{h}}\in\Omega^{p_{h}}$ and $\omega^{p_{l}}\in\Omega^{p_{l}}$, and the reward outcome sequences $\omega^{R_{h}}\in\Omega^{R_{h}}$ and $\omega^{R_{l}}\in\Omega^{R_{l}}$, are not shared between the high and low fidelity environments.
Formulating these separate outcome sets is conceptually helpful. However, practically, we only explicitly sample values for the policy outcomes $\omega_{0}^{\pi},\omega_{1}^{\pi},\ldots,\omega_{T}^{\pi}$ in our implementation. Algorithm 1 outlines the steps of the sampling process. First, the pre-sampled policy outcomes are used to rollout a high-fidelity trajectory $\tau^{h}$. Next, the initial state in the low-fidelity environment is fixed to match the initial state of $\tau^{h}$ and therefore to effectively enforce a common $\omega^{\Delta_{s_{I}}}$. Finally, the same sequence of pre-sampled policy outcomes is then used to generate the low-fidelity trajectory $\tau^{l}$.
@@
Correlated action sampling via policy distribution reparameterization. In order to use the sampled outcomes $\omega_{t}^{\pi}$ to deterministically select an action under the parameterized policy $\pi_{\theta}$, we implement a technique inspired by the so-called reparameterization trick used in variational autoencoders (Kingma, 2013). In particular, in continuous action spaces, we draw $\omega_{t}^{\pi}\sim\mathcal{N}(0,1)$, and the policy $\pi_{\theta}(s_{t},\omega_{t}^{\pi})$ is trained to output a state-dependent mean and standard deviation which are used to transform $\omega_{t}^{\pi}$ into an action $a_{t}$. Meanwhile, in discrete action spaces, one can draw $\omega^{\pi}_{t}\sim\textrm{Uniform}(0,1)$ and apply the Gumbel-Max trick (Huijben et al., 2022) to sample $a_{t}$ according to the state-dependent probability distribution defined by the policy $\pi_{\theta}(s_{t},\omega_{t}^{\pi})$.
@@
Defining multi-fidelity variant of the REINFORCE algorithm. To this point, we have defined a mechanism for sampling correlated trajectories $\tau^{h}$ and $\tau^{l}$ from multi-fidelity environments $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$, as well as a framework for using said trajectories to construct reduced-variance estimators of the policy gradient from trajectory-dependent variables $X_{\tau^{h}}^{\pi_{\theta}}$ and $X_{\tau^{l}}^{\pi_{\theta}}$. With these elements of the MFPG framework in place, we may implement a multi-fidelity variant of the REINFORCE algorithm. We note that, in principle, MFPG can also be instantiated with other on-policy policy gradient methods, by replacing the way in which the value of $X_{\tau}^{\pi_{\theta}}$ is computed from sampled trajectories. In this work, we focus on multi-fidelity REINFORCE in order to ease theoretical analysis, leaving extensions to other algorithms for future work.
@@
We implement a multi-fidelity variant of the REINFORCE algorithm (Williams, 1988; 1992) by defining $X_{\tau}^{\pi_{\theta}}\vcentcolon=\frac{1}{T}\sum_{t=0}^{T-1}G_{t}\log\pi_{\theta}(a_{t}|s_{t})$, as described above. Additionally, to match the commonly used variance-reduced variant of the REINFORCE algorithm (Peters & Schaal, 2006), we subtract state values estimated by a value network from the Monte Carlo returns, i.e., $X_{\tau}^{\pi_{\theta}}\vcentcolon=\frac{1}{T}\sum_{t=0}^{T-1}(G_{t}-V_{\phi}(s_{t}))\log\pi_{\theta}(a_{t}|s_{t})=\frac{1}{T}\sum_{t=0}^{T-1}A_{\phi}(s_{t},a_{t})\log\pi_{\theta}(a_{t}|s_{t})$. Here, $V_{\phi}(s_{t}),A_{\phi}(s_{t},a_{t})$ denote value and advantage functions estimated from samples. A common advantage network $A_{\phi}$ is used to compute both $X_{\tau^{h}}^{\pi_{\theta}}$ and $X_{\tau^{l}}^{\pi_{\theta}}$, and is trained using sampled tuples $(s^{h}_{t},a^{h}_{t},r_{t}^{h},s_{t+1}^{h})$ from the high-fidelity dataset. Algorithm˜2 in Appendix˜A summarizes the MFPG REINFORCE algorithm.
@@ 5 Theoretical Analysis of Multi-Fidelity Policy Gradients
Assumption 5 (High-Fidelity Policy Gradient Estimate).
The high-fidelity policy gradient estimate $\nabla_{\theta}X_{\tau^{h}}^{\pi_{\theta}}$ is unbiased and has bounded variance, i.e., $\mathbb{E}\left[\nabla_{\theta}X_{\tau^{h}}^{\pi_{\theta}}\right]=\nabla_{\theta}J(\theta)$, and there exists $\sigma>0$ such that $\mathbb{E}\left[\left\|\nabla_{\theta}X_{\tau^{h}}-\nabla_{\theta}J(\theta)\right\|^{2}\right]\leq\sigma^{2}$.
Assumption˜5 ensures that the high-fidelity policy gradient estimator is a “good” estimator, in the sense that the high-fidelity policy gradient would have learned a good policy if it had unrestricted access to high-fidelity data. Such an assumption is required to formally show convergence (see, e.g., Papini et al. (2018)).
Finally, in order for the proposed approach to be effective, there must be nonzero correlation between random variables computed from the high- and low-fidelity environments, i.e., the coefficient $c^{*}$ in Eq.˜1 is nonzero. Hence, we make the following assumption.
@@
Assumption˜6 can be interpreted as a comment on the “goodness" of the low-fidelity estimator: while the low-fidelity estimator may be inaccurate and have return values that are incomparable to the high-fidelity due to different transition dynamics in the two environments, the assumption ensures that the lower fidelity still preserves some meaningful connection with the higher fidelity. This connection is captured by the correlation between the two fidelities.
Under these assumptions, we establish that the variance-reduced MFPG iterates lead to a faster decrease in policy gradient norms in comparison to the high-fidelity-only iterates, when employed for a finite horizon REINFORCE algorithm. Further, we establish the global asymptotic convergence of the MFPG to a first-order stationary point of the high-fidelity objective function. The proof is presented in Appendix˜B.
@@
where $L_{T}$ is the Lipschitz constant of the high-fidelity policy gradient, established in Lemma˜1. Moreover, we recover the corresponding rate for the high-fidelity only policy iterates $\{\theta_{k}^{h}\}_{k\in\mathbb{Z}^{+}}$ by substituting $\rho=0$. Finally, we have that $\lim_{k\to\infty}\mathbb{E}\left[\|\nabla J(\theta_{k}^{MFPG})\|\right]=0$ almost surely.
@@
The non-asymptotic bounds leverage the Lipschitzness of the high-fidelity-only policy gradient estimator, which we establish in Lemma˜1 in Appendix˜B, while the asymptotic convergence proof utilizes the supermartingale convergence theorem (Robbins & Siegmund, 1971) as follows. We first show that the norm of the MFPG estimator is bounded, then construct a specific non-negative auxiliary random variable that we show is a supermartingale and apply the supermartingale theorem to establish the result. The complete proof is presented in Appendix˜B.
@@ 6 Experiments
This section empirically evaluates the proposed MFPG algorithm in a suite of simulated benchmark RL tasks. In high-fidelity data-scarce regimes, the experimental results demonstrate the following key insights:
Key insights: (I1) When the multi-fidelity dynamics gap is mild, MFPG can substantially reduce the variance of PG estimates compared to standard variance reduction via state-value baseline subtraction. (I2) With mild to moderate dynamics gaps, MFPG reliably outperforms the baseline that only uses high-fidelity data in median performance, matching the leading multi-fidelity methods despite its simplicity and minimal tuning overhead. In more challenging scenarios with large gaps, MFPG is the most robust among all evaluated multi-fidelity methods, incurring the fewest performance collapses. (I3) MFPG can remain effective even when low-fidelity samples exhibit reward misspecification.
@@ 6.1 Experiment Setup
Task settings. We evaluate our approach on the off-dynamics RL benchmark introduced by Lyu et al. (2024b), focusing on three standard MuJoCo tasks (Todorov et al., 2012): (i) Hopper-v3, (ii) Walker2d-v3, and (iii) HalfCheetah-v3.111The benchmark (Lyu et al., 2024b) also contains the Ant-v3 environment. Unlike the other three tasks, the observation space and reward function of Ant-v3 also depend on contact forces, which cannot be reset to specific values by the user. Because of the difficulty in matching initial states across fidelities, we focus our evaluations on the other three tasks. In each task, the low-fidelity environment corresponds to the original dynamics, while the high-fidelity variants include changes in gravity, friction, and kinematics. Gravity and friction are varied across five levels $\{0.5\times,0.8\times,1.2\times,2.0\times,5.0\times\}$; the intermediate $0.8\times$ and $1.2\times$ levels are not part of the original benchmark but are introduced here to examine settings with mild dynamics shifts. Kinematic variations are categorized as easy, medium, or hard, each constraining a specific joint to a reduced range of motion to mimic actuator damage. For kinematic shifts, we consider cases where the agent’s thigh joint is impaired (leg joint for Hopper-v3, since the original benchmark does not include a thigh joint shift for Hopper). In total, the task set spans 39 distinct settings.
@@
Baselines. We compare MFPG against the following baselines: High-Fidelity Only, which trains solely on the limited high-fidelity samples; Low-Fidelity Only (100×), which trains on abundant low-fidelity data (100× more samples) and is evaluated in the high-fidelity environment; DARC (Eysenbach et al., ), which learns domain classifiers to estimate dynamics gaps for the sampled low-fidelity transitions and uses the estimated gaps to augment low-fidelity rewards; PAR (Lyu et al., 2024a), which augments low-fidelity rewards using dynamics mismatch estimated in a learned latent space; and More High-Fidelity Data (15×), an oracle baseline that trains with access to 15× more high-fidelity samples. Please refer to Appendix˜D for a more detailed description of the baselines.
@@
We instantiate the MFPG framework with the on-policy policy gradient algorithm REINFORCE (Williams, 1992) with state value subtraction. DARC (Eysenbach et al., ) and PAR (Lyu et al., 2024a) were originally demonstrated with the off-policy soft actor-critic algorithm (Haarnoja et al., 2018). To isolate the effect of their multi-fidelity mechanisms from the backbone algorithm and enable fair comparison, we adapt both PAR and DARC to the REINFORCE backbone. Our code is available at https://github.com/CLeARoboticsLab/MultiFidelityPolicyGradients.
@@
Experiment regime. We evaluate MFPG in a regime where high-fidelity samples are scarce, reflecting many real-world applications such as robot learning and molecule design. To simulate this setting, we restrict all methods—except the oracle baseline More High-Fidelity Data (15×)—to a buffer of 100 high-fidelity samples per policy update. Multi-fidelity methods may supplement this limited buffer with up to 100× additional low-fidelity samples per update. Training for all methods is conducted for up to 1M high-fidelity environment steps, except for the More High-Fidelity Data (15×) baseline, which is trained for 15M high-fidelity steps.
@@
Hyperparameter configuration. Given the restriction on high-fidelity samples per policy update, we first tune hyperparameters for the High-Fidelity Only REINFORCE baseline (with state-value baselines subtracted). This tuned configuration serves as the common backbone for all multi-fidelity methods (MFPG, DARC, and PAR), in line with standard practice in the off-dynamics RL literature (Xu et al., 2023; Lyu et al., 2024a). Their method-specific hyperparameters, including the number of low-fidelity samples for DARC and PAR, are tuned separately. The Low-Fidelity Only (100×) baseline is implemented as a direct ablation of the multi-fidelity methods. Finally, the oracle baseline More High-Fidelity Data (15×) is additionally tuned to ensure fair benefit from the extra high-fidelity samples. For further details on hyperparameter setup, see Appendix˜D. Overall, our tuning protocol grants multi-fidelity baselines more tuning than MFPG, underscoring the simplicity and minimal tuning overhead of MFPG.
@@
Notably, in our tuning process, PAR is particularly sensitive to hyperparameters; cf. Fig.˜14 in Appendix˜D. Following (Lyu et al., 2024a), we adopt separate configurations per task for PAR, whereas all other methods use a single configuration across settings. MFPG has the fewest hyperparameters among all multi-fidelity methods. We tuned only the exponential moving-average weight $\eta_{\text{ma}}$ in Eq.˜2 and found the performance to be insensitive to this choice.
Result reporting. We evaluate all methods using five random seeds. Since five-seed results remain noisy and standard deviations are less reliable with a smaller number of runs, we report median, maximum, and minimum values. In bar plots, bar heights denote medians and error bars indicate maxima and minima; in line plots, solid lines represent medians and shaded regions correspond to maxima/minima.
@@
Performance is evaluated in the high-fidelity environment every 2000 high-fidelity steps, using 10 evaluation episodes with a deterministic policy each time. We adopt two primary metrics: (i) the final evaluation return, defined as the median/max/min return averaged over the last 20 evaluation steps, and (ii) area under curve (AUC), the area under the median/max/min evaluation-return curve across training, estimated by integrating evaluation return along the high-fidelity step axis using the composite trapezoidal rule. AUC captures a method’s accumulated return performance (Stadie et al., 2015; Osband et al., 2016; Hessel et al., 2018). Finally, to assess robustness, we also report the number of performance collapses, defined as cases where the median of a method falls below 50% of the median value of the corresponding High-Fidelity Only baseline.

### v4

@@ 20 $\tau^{l}_{i}\leftarrow s_{0}^{l},a_{0}^{l},r_{0}^{l},\ldots,s_{T}^{l}$
Correlated action sampling via policy distribution reparameterization.
@@
In order to use the sampled outcomes $\omega_{t}^{\pi}$ to deterministically select an action under the parameterized policy $\pi_{\theta}$, we implement a technique inspired by the so-called reparameterization trick used in variational autoencoders (Kingma, 2013). In particular, in continuous action spaces, we draw $\omega_{t}^{\pi}\sim\mathcal{N}(0,1)$, and the policy $\pi_{\theta}(s_{t},\omega_{t}^{\pi})$ is trained to output a state-dependent mean and standard deviation which are used to transform $\omega_{t}^{\pi}$ into an action $a_{t}$. Meanwhile, in discrete action spaces, one can draw $\omega^{\pi}_{t}\sim\textrm{Uniform}(0,1)$ and apply the Gumbel-Max trick (Huijben et al., 2022) to sample $a_{t}$ according to the state-dependent probability distribution defined by the policy $\pi_{\theta}(s_{t},\omega_{t}^{\pi})$.
To summarize, Algorithm 1 outlines the sampling procedure. First, we use the pre-sampled policy action-noise sequence to roll out a high-fidelity trajectory $\tau^{h}$. Next, the initial state in the low-fidelity environment is fixed to match the initial state of $\tau^{h}$ and thereby effectively enforcing a common $\omega^{\Delta_{s_{I}}}$. Finally, we reuse the same sequence of action-noise samples to generate the low-fidelity trajectory $\tau^{l}$.
Defining multi-fidelity variant of the REINFORCE algorithm.
@@
To this point, we have defined a mechanism for sampling correlated trajectories $\tau^{h}$ and $\tau^{l}$ from multi-fidelity environments $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$, as well as a framework for using said trajectories to construct reduced-variance estimators of the policy gradient from trajectory-dependent variables $X_{\tau^{h}}^{\pi_{\theta}}$ and $X_{\tau^{l}}^{\pi_{\theta}}$. With these elements of the MFPG framework in place, we may implement a multi-fidelity variant of the REINFORCE algorithm. We note that, in principle, MFPG can also be instantiated with other on-policy policy gradient methods, by replacing the way in which the value of $X_{\tau}^{\pi_{\theta}}$ is computed from sampled trajectories. In this work, we focus on multi-fidelity REINFORCE in order to ease theoretical analysis, leaving extensions to other algorithms for future work.
@@
We implement a multi-fidelity variant of the REINFORCE algorithm (Williams, 1988; 1992) by defining $X_{\tau}^{\pi_{\theta}}\vcentcolon=\frac{1}{T}\sum_{t=0}^{T-1}G_{t}\log\pi_{\theta}(a_{t}|s_{t})$, as described above. Additionally, to match the commonly used variance-reduced variant of the REINFORCE algorithm (Peters & Schaal, 2006), we subtract state values estimated by a value network from the Monte Carlo returns, i.e., $X_{\tau}^{\pi_{\theta}}\vcentcolon=\frac{1}{T}\sum_{t=0}^{T-1}(G_{t}-V_{\phi}(s_{t}))\log\pi_{\theta}(a_{t}|s_{t})=\frac{1}{T}\sum_{t=0}^{T-1}A_{\phi}(s_{t},a_{t})\log\pi_{\theta}(a_{t}|s_{t})$. Here, $V_{\phi}(s_{t}),A_{\phi}(s_{t},a_{t})$ denote value and advantage functions estimated from samples. We sample trajectories from the high- and low-fidelity environments to compute Monte Carlo returns $G^{h}_{t}$ and $G^{l}_{t}$, respectively. We then use a shared value function $V_{\phi}$ learned from high-fidelity samples $(s^{h}_{t},a^{h}_{t},r_{t}^{h},s_{t+1}^{h})$, to compute state-value baselines that are subtracted from both $G^{h}_{t}$ and $G^{l}_{t}$. This is an implementation choice that simplifies training; moreover, subtracting any state-dependent baseline does not change the expectation of the REINFORCE gradient estimator (Williams, 1992). Alternatively, one can train separate value functions for the high- and low-fidelity Monte Carlo returns. Algorithm˜2 in Appendix˜A summarizes the MFPG REINFORCE algorithm.
@@ 5 Theoretical Analysis of Multi-Fidelity Policy Gradients
Assumption 5 (High-Fidelity Policy Gradient Estimator).
The high-fidelity policy gradient estimator $\nabla_{\theta}X_{\tau^{h}}^{\pi_{\theta}}$ is unbiased and has bounded variance, i.e., $\mathbb{E}\left[\nabla_{\theta}X_{\tau^{h}}^{\pi_{\theta}}\right]=\nabla_{\theta}J(\theta)$, and there exists $\sigma>0$ such that $\mathbb{E}\left[\left\|\nabla_{\theta}X_{\tau^{h}}-\nabla_{\theta}J(\theta)\right\|^{2}\right]\leq\sigma^{2}$.
Assumption˜5 ensures that the high-fidelity policy gradient estimator is a “good” estimator, in the sense that the high-fidelity policy gradient would have learned a good policy if it had unrestricted access to high-fidelity data. Such an assumption is required to formally show convergence (see, e.g., Papini et al. (2018)). We note that Assumption˜5 is an assumption on the underlying high-fidelity policy gradient estimator; we do not require any policy gradient estimates computed from finite samples to be exact. In particular, the classic Monte Carlo REINFORCE policy gradient estimator considered in this work satisfies this requirement (Sutton et al., 1999). Extending our analysis to actor–critic-style estimators is nontrivial and is discussed as a limitation and direction for future work in Appendix˜G.
Finally, in order for the proposed approach to be effective, there must be nonzero correlation between random variables computed from the high- and low-fidelity environments, i.e., the coefficient $c^{*}$ in Eq.˜2 is nonzero. Hence, we make the following assumption.
@@
Assumption˜6 can be interpreted as a comment on the quality of the low-fidelity estimator: while the low-fidelity estimator may be inaccurate and have return values that are incomparable to the high-fidelity due to different transition dynamics in the two environments, the assumption ensures that the lower fidelity still preserves some meaningful connection with the higher fidelity. This connection is captured by the correlation between the two fidelities.
Under these assumptions, we establish that the variance-reduced MFPG iterates lead to a faster decrease in policy gradient norms in comparison to the high-fidelity-only iterates, when employed for a finite horizon REINFORCE algorithm. Further, we establish the global asymptotic convergence of the MFPG to a first-order stationary point of the high-fidelity objective function. The proof is presented in Appendix˜C.
@@
where $L_{T}$ is the Lipschitz constant of the high-fidelity policy gradient, established in Lemma˜2. Moreover, we recover the corresponding rate for the high-fidelity only policy iterates $\{\theta_{k}^{h}\}_{k\in\mathbb{Z}^{+}}$ by substituting $\rho=0$. Finally, we have that $\lim_{k\to\infty}\mathbb{E}\left[\|\nabla J(\theta_{k}^{MFPG})\|\right]=0$ almost surely.
@@
The non-asymptotic bounds leverage the Lipschitzness of the high-fidelity-only policy gradient estimator, which we establish in Lemma˜2 in Appendix˜C, while the asymptotic convergence proof utilizes the supermartingale convergence theorem (Robbins & Siegmund, 1971) as follows. We first show that the norm of the MFPG estimator is bounded, then construct a specific non-negative auxiliary random variable that we show is a supermartingale and apply the supermartingale theorem to establish the result. The complete proof is presented in Appendix˜C.
@@ 6 Experiments
This section empirically evaluates the proposed MFPG algorithm in a suite of simulated benchmark RL tasks. In high-fidelity data-scarce regimes, the experimental results demonstrate the following key insights [^1] Code: https://xinjie-liu.github.io/mfpg-rl/:
  • (I1, Variance reduction) When the multi-fidelity dynamics gap is mild, MFPG can substantially reduce the variance of PG estimates compared to standard variance reduction via state-value baseline subtraction.
  • (I2, Reliability and robustness) MFPG provides a reliable and robust way to exploit low-fidelity data. In our baseline comparisons in Section˜6.3, for scenarios where low-fidelity data are neutral or beneficial (Hopper-gravity, Walker2d-friction) and dynamics gaps are mild to moderate (0.5× to 2.0×), MFPG is, among the evaluated off-dynamics RL and low-fidelity-only approaches, the only method that consistently and statistically significantly improves the mean performance over a baseline trained solely on high-fidelity data in 8 out of 8 scenarios. When low-fidelity data become harmful, MFPG exhibits the strongest robustness against performance degradation among the evaluated off-dynamics RL and low-fidelity-only methods. Strong off-dynamics RL methods tend to exploit low-fidelity data more aggressively, and can achieve higher performance than MFPG at times (e.g., in Walker2d) but can also fail substantially more severely (e.g., in HalfCheetah).
  • (I3, Misspecified reward) MFPG can remain effective even when low-fidelity samples exhibit reward misspecification, e.g., when the low-fidelity reward is the negative (anti-correlated) version of the high-fidelity reward.
@@ 6.1 Experiment Setup
We evaluate our approach on the off-dynamics RL benchmark introduced by Lyu et al. (2024b), focusing on three standard MuJoCo tasks (Todorov et al., 2012): (i) Hopper-v3, (ii) Walker2d-v3, and (iii) HalfCheetah-v3. In each task, the low-fidelity environment corresponds to the original dynamics, while the high-fidelity variants include changes in gravity or friction. Gravity and friction are varied across five levels $\{0.5\times,0.8\times,1.2\times,2.0\times,5.0\times\}$; the intermediate $0.8\times$ and $1.2\times$ levels are not part of the original benchmark but are introduced here to examine settings with mild dynamics shifts. In total, the task set spans 15 distinct settings. In Sections˜6.3, 6.4 and 6.5, we evaluate each setting with 20 random seeds per method.
We provide additional results on 24 further task settings in Appendix˜E, evaluated with 5 random seeds per method. Due to the small number of seeds, these additional results are primarily used to illustrate the consistency of qualitative trends, and we do not draw formal statistical conclusions from them.
We note that the benchmark (Lyu et al., 2024b) also includes the Ant-v3 environment. Unlike the other three tasks, the observation space and reward function in Ant-v3 depend explicitly on contact forces, which cannot be reliably reset to user-specified values, as required by our MFPG sampling scheme. Due to this difficulty in matching initial states across fidelities, we focus our evaluations on the other three tasks described above.
Baselines.
@@
We compare MFPG against the following baselines: High-Fidelity Only, which trains solely on the limited high-fidelity samples; Low-Fidelity Only (100×), which trains on abundant low-fidelity data (100× more samples) and is evaluated in the high-fidelity environment; DARC (Eysenbach et al., ), which learns domain classifiers to estimate dynamics gaps for the sampled low-fidelity transitions and uses the estimated gaps to augment low-fidelity rewards; and PAR (Lyu et al., 2024a), which augments low-fidelity rewards using dynamics mismatch estimated in a learned latent space. Please refer to Appendix˜F for a more detailed description of the baselines.
@@
We instantiate the MFPG framework with the on-policy policy gradient algorithm REINFORCE (Williams, 1992) with state value subtraction. DARC (Eysenbach et al., ) and PAR (Lyu et al., 2024a) were originally demonstrated with the off-policy soft actor-critic algorithm (Haarnoja et al., 2018). To isolate the effect of their multi-fidelity mechanisms from the backbone algorithm and enable fair comparison, we adapt both PAR and DARC to the REINFORCE backbone. We include our code in the supplementary materials and will release the project as open source upon acceptance.
Experiment regime.
@@
We evaluate MFPG in a regime where high-fidelity samples are scarce, reflecting many real-world applications such as robot learning and molecule design. To simulate this setting, we restrict all methods to a buffer of 100 high-fidelity samples per policy update. Multi-fidelity methods may supplement this limited buffer with up to 100× additional low-fidelity samples per update. Training for all methods is conducted for up to 1M high-fidelity environment steps.
Hyperparameter configuration.
@@
Given the restriction on high-fidelity samples per policy update, we first tune hyperparameters for the High-Fidelity Only REINFORCE baseline (with state-value baselines subtracted). This tuned configuration serves as the common backbone for all multi-fidelity methods (MFPG, DARC, and PAR), in line with standard practice in the off-dynamics RL literature (Xu et al., 2023; Lyu et al., 2024a). Their method-specific hyperparameters, including the number of low-fidelity samples for DARC and PAR, are tuned separately. The Low-Fidelity Only (100×) baseline is implemented as a direct ablation of the multi-fidelity methods. For further details on hyperparameter setup, see Appendix˜F. Overall, our tuning protocol grants the tested multi-fidelity baselines more tuning than MFPG, underscoring the simplicity and minimal tuning overhead of MFPG.
@@
Notably, in our tuning process, PAR is particularly sensitive to hyperparameters; cf. Fig.˜18 in Appendix˜F. Following (Lyu et al., 2024a), we adopt separate configurations per task for PAR, whereas all other methods use a single configuration across settings. MFPG has the fewest hyperparameters among all the tested multi-fidelity methods. We tuned only the exponential moving-average weight $\eta_{\text{ma}}$ in Eq.˜3 and found that MFPG ’s median performance consistently exceeded that of the High-Fidelity Only baseline across the tested values of $\eta_{\text{ma}}$, even though the absolute performance metrics varied with $\eta_{\text{ma}}$.
Result reporting.
For the variance reduction study in Section˜6.2, we run $5$ random seeds per method, as this experiment already aggregates a large number of gradient estimates within each run and is primarily intended to illustrate variance trends. For the experiments in Sections˜6.3, 6.4 and 6.5, we run $20$ random seeds per method and task setting.
@@
Performance is evaluated in the high-fidelity environment every 2000 high-fidelity steps, using 10 evaluation episodes with a deterministic policy each time. We adopt two primary metrics: (i) the final evaluation return, defined as the return averaged over the last 20 evaluation steps, and (ii) area under curve (AUC), the area under the evaluation-return curve across training, estimated by integrating evaluation return along the high-fidelity step axis using the composite trapezoidal rule. AUC captures a method’s accumulated return performance (Stadie et al., 2015; Osband et al., 2016; Hessel et al., 2018). Finally, to assess robustness, we also report the number of performance collapses, defined as cases where the median of a method falls below 50% of the median value of the corresponding High-Fidelity Only baseline.
Uncertainty and significance.
Where reported, we compute two-sided $95\%$ nonparametric bootstrap confidence intervals using R=10, [^0] R=10{,}000 resamples (sampling with replacement over seeds). We report point estimates as sample means. In the baseline comparisons in Section˜6.3, we bootstrap the mean difference of each method compared to High-Fidelity Only, i.e., $\Delta=\text{mean}_{\text{method}}-\text{mean}_{\text{High-Fidelity Only}}$, by independently resampling seeds within each method and recomputing $\Delta$ for each resample. We declare an improvement statistically significant at $\alpha=0.05$ when the $95\%$ bootstrap confidence interval for $\Delta$ excludes zero.